Find the y-intercept -x+3y=15
A). (5,0)

B) (0,-15)
C) (0,5)
D) (-15,0)

Answer by checkley71(8403)

-x+3y=15 or 3y=x+15 or y=x/3+5 or the y intercept is C)(0,5) because in the form Y=mX+b b=y intercept.